Evaluate the function.
Given that f(x) = (1)/(4-8x) , find f(r - 3).
A) (1)/(1-8r)
B) (1)/(28-8r)
C) (1)/(28+b)
D) (1)/(r-3-8x)

Answer by RAY100(1637)

f(x) = 1/(4-8x)
.
f(r-3) = 1/ { 4-8(r-3) }
.
f(r-3 = 1/ (4-8r+24)
.
f(r-3) = 1/ (28 -8r),,,,,,or B